From GlueXWiki
Jump to: navigation, search

March 22, 2012 FDC meeting


  1. Oxygen contamination
    • Test results FDC ELOG (Beni, Lubomir)
    • Future tests (Dave, Lubomir)
  2. Production Construction Tracking (Dave)
    • Production status
    • Testing at Blue Crab (Lubomir)
  3. Engineering update (Dave, Lubomir)
  4. Electronics update (Chris)
  5. Other


Participants: Tim, Dave, Simon, Beni, Chris, John, Casey, and Lubomir.


- Dave: Remaining to be done on package #3: last two wire frames to be electroplated and Chris has to put the elements on one of them; daughter cards on one type-2 cathode to be installed. Package #4: all wire frames were laminated, on three the caps were already replaced and tested and Anatoly is working on the fourth one; one wire frame strung. The plan is to stop the installation of the third package, stop the wire frame production after finishing the last one and stop at least by Monday the cathode production. The hope is by that time we will have an idea how to fix the oxygen leakage into the chambers.

- Lubomir: fourth cell of the third package was tested with a source and a scope. All channels are OK, but a leakage was found at the top end window.

Oxygen contamination in the first package

- Beni: after replacing the plastic rods on the first package with aluminum ones the pressure inside the chamber increased from 70 to 110 Pa. To keep the pressure the same we had to decrease the flow from 200 to 300 scc/min. So, the main effect was that we improved the tightness of the chamber significantly but the oxygen percentage improved just a little (~10%).

- Replacing the old O-rings (EPDM) with new (Buna-N, or Nitrile) on the testing chamber resulted in decreasing the oxygen from 1400 vpm to 940 vpm (in three days, last value stable for several hours) which is not enough. The configuration of the testing chamber was from bottom to top: Lexan, spacer ring with two new O-rings, cathode type-2, end-window with new O-ring.

- This morning we started a new test without the cathode and end-window so that we can get the results faster without the air-pockets in these elements. Configuration: Lexan and spacer ring from before (with the two new O-rings), empty type-1 cathode frame with an old type O-ring on the top, another Lexan sheet. The O-ring on the cathode frame was not glued with super glue (except at two points). If the main oxygen leakage is through the G10 groove/old O-ring we had similar O-ring configuration before but with cathode type-3 and end-window, and expect to see oxygen at ~1400vpm.

- Future plans: The idea is tomorrow to replace the cathode frame with another one on which the groove was coated today with Hysol epoxy. Dave and Casey made different sample coatings, also diluting the Hysol with ethanol, but we decided to use the non-diluted epoxy since it covers better the G10 fibers at the same time increasing the g10 thickness by only ~1mil. If coating the grooves results in significant oxygen reduction we may want to try next week to see also the effect of the coating AND the new type O-ring. Then we will plan the refurbishing procedures, will start with the third package. We will bring the oxygen sensor to Blue Crab and will measure the oxygen cell by cell before and after refurbishment.


- We discussed if the Fluorinert will be a problem for our chambers. Lubomir talked to experts and did some study in the literature but no evidence found either for or against the use of Fluorinert. Tim: it's not a problem to use water. Beni asked if the pressure must be higher with water. Tim explained: it depends on the temperature tolerance and if acceptable we can always lower the pressure. Still the thick tube material will be there. Using air for cooling is probably not an option.

- Dave: we have received the parts for the hydrostatic tests of the cooling system.


- Beni asked when the checking of all the pre-amps will be finished; he is interested actually in the fADC125 that is used in the tests. Beni's idea is to use at least the two available fADC125 to test the cathode channels on the first package; non of them is connected now.

- Chris: in principle this can be done within few weeks. We will discuss this with Fernando.